OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~that~\log _{2}x-\log _{x}8=2,find~the~values~of~x.\]

ganeshie8 (ganeshie8):

Hint : \[\log_y x = \dfrac{1}{\log_x y}\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ \log _{2}x }\]

ganeshie8 (ganeshie8):

\[\log_2 x - \log_x 8 = 2\] \[\log_2 x - \log_x 2^3 = 2\] \[\log_2 x - 3\log_x 2 = 2\] \[\log_2 x - 3\dfrac{1}{\log_2 x} = 2\]

ganeshie8 (ganeshie8):

let \(t = \log_2 x\) : \[t - 3\frac{1}{t} = 2\]

ganeshie8 (ganeshie8):

turn it into a quadratic and solve \(t\)

OpenStudy (anonymous):

\[\log _{2}x(1-3)=2\]

ganeshie8 (ganeshie8):

nope \[t - 3\frac{1}{t} = 2 \] multiply \(t\) throught out : \[t^2 - 3 = 2t \] which is same as \[t^2-2t - 3 = 0 \] see if you can factor this and solve \(t\)

OpenStudy (anonymous):



OpenStudy (anonymous):

t=3 and t=-2

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

\[\log _{2}x(\log _{2}x-3\frac{ 1 }{ \log _{2}x })=2\]

ganeshie8 (ganeshie8):

heyy first solve the quadratic

ganeshie8 (ganeshie8):

t=3 and t=-2 are wrong

OpenStudy (anonymous):

\[t^2-2t-3=0\]\[t(t-2)-3=0\]\[t=3~and~t=2\]

ganeshie8 (ganeshie8):

\[t^2 - 2t -3 = 0\] \[t^2 - 3t+t -3 = 0\] \[t(t - 3)+1(t -3) = 0\] \[(t - 3)(t+1) = 0\] \[t = 3, -1\]

OpenStudy (anonymous):

okay :)

ganeshie8 (ganeshie8):

and previously we let \(\log_2 x=t \) so \[\log_2 x = 3 \implies x = 2^3\] \[\log_2 x = -1 \implies x = \frac{1}{2}\]

OpenStudy (anonymous):

x=8,x=1/2

OpenStudy (anonymous):

Thnx @ganeshie8 :)

ganeshie8 (ganeshie8):

yw :)